# Efficient Epistemic Uncertainty Estimation in Regression Ensemble Models Using Pairwise-Distance Estimators

### v3

@@ Appendix E Probabilistic Network Ensembles
Env Acq. Batch Random BatchBALD BADGE BAIT LCMD Monte Carlo KL Bhatt
hetero 10 0.13 $\pm$ 0.57 -0.67 $\pm$ 0.55 -0.72 $\pm$ 0.61 -0.78 $\pm$ 0.59 -0.46 $\pm$ 0.49 0.89 $\pm$ 0.1 0.87 $\pm$ 0.15 0.74 $\pm$ 0.3
25 0.03 $\pm$ 0.47 0.33 $\pm$ 0.76 0.38 $\pm$ 0.76 0.13 $\pm$ 0.82 0.41 $\pm$ 0.75 1.02 $\pm$ 0.11 0.95 $\pm$ 0.14 0.92 $\pm$ 0.24
50 0.23 $\pm$ 0.57 0.48 $\pm$ 0.71 0.14 $\pm$ 0.77 0.11 $\pm$ 0.85 0.17 $\pm$ 0.69 1.01 $\pm$ 0.15 1.1 $\pm$ 0.11 1.09 $\pm$ 0.12
100 0.23 $\pm$ 0.42 0.23 $\pm$ 0.78 0.28 $\pm$ 0.85 -0.03 $\pm$ 0.74 0.46 $\pm$ 0.7 1.1 $\pm$ 0.11 1.11 $\pm$ 0.17 1.08 $\pm$ 0.2
bimodal 10 -17.71 $\pm$ 10.01 0.32 $\pm$ 0.33 0.41 $\pm$ 0.21 -1.36 $\pm$ 4.14 0.2 $\pm$ 0.49 0.6 $\pm$ 0.1 0.59 $\pm$ 0.14 0.61 $\pm$ 0.07
25 -20.78 $\pm$ 5.64 0.61 $\pm$ 0.19 0.48 $\pm$ 0.23 0.57 $\pm$ 0.21 0.57 $\pm$ 0.21 0.66 $\pm$ 0.03 0.66 $\pm$ 0.03 0.67 $\pm$ 0.04
50 -12.93 $\pm$ 8.55 0.66 $\pm$ 0.09 0.59 $\pm$ 0.16 0.61 $\pm$ 0.18 0.67 $\pm$ 0.09 0.69 $\pm$ 0.02 0.69 $\pm$ 0.03 0.69 $\pm$ 0.03
100 -5.79 $\pm$ 7.19 0.66 $\pm$ 0.1 0.64 $\pm$ 0.11 0.68 $\pm$ 0.08 0.67 $\pm$ 0.1 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02
Pendulum 10 2.35 $\pm$ 2.94 -10.76 $\pm$ 7.28 -15.25 $\pm$ 8.86 -9.76 $\pm$ 9.09 -3.4 $\pm$ 2.6 4.94 $\pm$ 0.48 5.51 $\pm$ 0.72 5.21 $\pm$ 0.61
25 4.19 $\pm$ 0.89 -0.4 $\pm$ 14.08 0.2 $\pm$ 13.03 4.68 $\pm$ 2.98 5.42 $\pm$ 5.25 5.45 $\pm$ 0.43 6.21 $\pm$ 0.98 6.23 $\pm$ 0.84
50 4.34 $\pm$ 0.75 8.34 $\pm$ 0.83 7.97 $\pm$ 0.74 8.68 $\pm$ 0.72 8.13 $\pm$ 0.94 5.92 $\pm$ 0.32 7.65 $\pm$ 0.5 7.15 $\pm$ 0.76
100 4.69 $\pm$ 0.56 9.53 $\pm$ 0.96 7.07 $\pm$ 5.49 8.8 $\pm$ 1.23 9.27 $\pm$ 0.83 6.09 $\pm$ 0.42 7.66 $\pm$ 0.64 7.77 $\pm$ 0.68
Hopper 10 -31.94 $\pm$ 28.24 -26.65 $\pm$ 15.57 -12.96 $\pm$ 9.55 -10.81 $\pm$ 4.25 -2.04 $\pm$ 5.45 9.36 $\pm$ 1.79 8.55 $\pm$ 1.34 8.55 $\pm$ 1.88
25 -5.74 $\pm$ 2.66 12.62 $\pm$ 4.32 13.88 $\pm$ 2.27 17.29 $\pm$ 3.45 16.69 $\pm$ 1.7 13.31 $\pm$ 2.3 13.9 $\pm$ 1.3 13.93 $\pm$ 1.6
50 -2.31 $\pm$ 3.08 27.28 $\pm$ 2.42 23.49 $\pm$ 2.97 28.08 $\pm$ 2.76 24.73 $\pm$ 1.71 16.43 $\pm$ 1.27 17.6 $\pm$ 1.89 16.93 $\pm$ 1.43
100 1.75 $\pm$ 2.83 30.35 $\pm$ 1.44 26.54 $\pm$ 2.19 28.44 $\pm$ 1.52 26.22 $\pm$ 2.75 18.5 $\pm$ 0.65 20.19 $\pm$ 1.37 18.54 $\pm$ 1.23
Ant 10 11.83 $\pm$ 2.29 -4.91 $\pm$ 12.37 3.29 $\pm$ 6.43 10.2 $\pm$ 5.57 12.39 $\pm$ 3.69 13.8 $\pm$ 5.87 22.13 $\pm$ 3.11 23.32 $\pm$ 2.73
25 17.78 $\pm$ 2.28 5.79 $\pm$ 7.68 8.44 $\pm$ 3.68 11.97 $\pm$ 5.41 12.37 $\pm$ 4.35 18.69 $\pm$ 9.06 26.98 $\pm$ 1.64 27.66 $\pm$ 1.78
50 22.21 $\pm$ 2.11 12.24 $\pm$ 3.25 14.94 $\pm$ 3.3 16.16 $\pm$ 3.2 14.41 $\pm$ 3.79 20.54 $\pm$ 8.88 29.84 $\pm$ 1.09 29.62 $\pm$ 0.7
100 23.4 $\pm$ 1.85 15.26 $\pm$ 6.5 16.87 $\pm$ 3.75 19.6 $\pm$ 2.07 16.08 $\pm$ 2.67 20.03 $\pm$ 9.87 31.24 $\pm$ 1.09 31.63 $\pm$ 0.9
Humanoid 10 -230.41 $\pm$ 19.87 -156.59 $\pm$ 82.79 -210.52 $\pm$ 53.23 -242.95 $\pm$ 2.96 -244.1 $\pm$ 3.79 -247.83 $\pm$ 1.71 -226.42 $\pm$ 46.74 -194.78 $\pm$ 73.74
25 -245.45 $\pm$ 2.24 -224.95 $\pm$ 37.12 -245.82 $\pm$ 5.03 -247.3 $\pm$ 1.26 -245.87 $\pm$ 4.85 -247.83 $\pm$ 1.19 -247.39 $\pm$ 3.18 -232.34 $\pm$ 42.02
50 -246.91 $\pm$ 1.28 -247.38 $\pm$ 3.28 -249.28 $\pm$ 0.52 -249.2 $\pm$ 0.52 -248.59 $\pm$ 1.68 -248.06 $\pm$ 1.82 -249.41 $\pm$ 0.73 -223.49 $\pm$ 77.83
100 -247.74 $\pm$ 0.6 -248.97 $\pm$ 1.1 -249.4 $\pm$ 0.6 -249.66 $\pm$ 0.53 -248.13 $\pm$ 1.25 -247.25 $\pm$ 3.36 -249.91 $\pm$ 0.12 -223.29 $\pm$ 79.8
@@ Appendix F Introduction to Normalizing Flows
NFs are powerful non-parametric models that have demonstrated the ability to fit flexible multi-modal distributions (Tabak & Vanden-Eijnden, 2010; Tabak & Turner, 2013). These models achieve this by transforming a simple base continuous distribution, such as Gaussian or Beta, into a more complex one using the change of variable formula. By enabling scoring and sampling from the fitted distribution, NFs find application across various problem domains. Let $B$ represent the base distribution, a D-dimensional continuous random vector with $p_{B}(b)$ as its density function, and let $Y=g(B)$, where $g$ is an invertible function with an existing inverse $g^{-1}$, and both $g$ and $g^{-1}$ are differentiable. Leveraging the change of variable formula, we can express the distribution of $Y$ as follows:
@@
where $J(\cdot)$ denotes the Jacobian, and $\det$ signifies the determinant. The first term on the right-hand side of Equation 12 governs the shape of the distribution, while $|\det(J(g^{-1}(y)))|$ normalizes it, ensuring the distribution integrates to one. Complex distributions can be effectively modeled by making $g(b)$ a learnable function with tunable parameters $\theta$, denoted as $g_{\theta}(b)$. However, it is essential to select $g$ carefully to guarantee its invertibility and differentiability. For examples of suitable choices, please refer to (Papamakarios et al., 2021).
@@ Appendix G Hypothesis Testing Details
We conducted Welch’s t-tests to compare means $(\mu_{i},\mu_{j})$ between different estimators, as this test relaxes the assumption of equal variances compared to other hypothesis tests (Colas et al., 2019). The means for both KL and Bhatt were compared to each of the baseline methods: BatchBALD, BADGE, BAIT, LCMD, MC and random. To control the family-wise error rate (FWER), we performed a Holm-Bonferroni correction across each setting, environment, and acquisition batch. This follows best practices to ensure our results are statistically significant and do not occur just by random chance.

### v4

@@ Appendix E Probabilistic Network Ensembles
Env Acq. Batch Random BatchBALD BADGE BAIT MC (BALD) KL (ours) Bhatt (ours)
hetero 10 0.13 $\pm$ 0.57 -0.67 $\pm$ 0.55 -0.72 $\pm$ 0.61 -0.78 $\pm$ 0.59 0.89 $\pm$ 0.1 0.87 $\pm$ 0.15 0.74 $\pm$ 0.3
25 0.03 $\pm$ 0.47 0.33 $\pm$ 0.76 0.38 $\pm$ 0.76 0.13 $\pm$ 0.82 1.02 $\pm$ 0.11 0.95 $\pm$ 0.14 0.92 $\pm$ 0.24
50 0.23 $\pm$ 0.57 0.48 $\pm$ 0.71 0.14 $\pm$ 0.77 0.11 $\pm$ 0.85 1.01 $\pm$ 0.15 1.1 $\pm$ 0.11 1.09 $\pm$ 0.12
100 0.23 $\pm$ 0.42 0.23 $\pm$ 0.78 0.28 $\pm$ 0.85 -0.03 $\pm$ 0.74 1.1 $\pm$ 0.11 1.11 $\pm$ 0.17 1.08 $\pm$ 0.2
bimodal 10 -17.71 $\pm$ 10.01 0.32 $\pm$ 0.33 0.41 $\pm$ 0.21 -1.36 $\pm$ 4.14 0.6 $\pm$ 0.1 0.59 $\pm$ 0.14 0.61 $\pm$ 0.07
25 -20.78 $\pm$ 5.64 0.61 $\pm$ 0.19 0.48 $\pm$ 0.23 0.57 $\pm$ 0.21 0.66 $\pm$ 0.03 0.66 $\pm$ 0.03 0.67 $\pm$ 0.04
50 -12.93 $\pm$ 8.55 0.66 $\pm$ 0.09 0.59 $\pm$ 0.16 0.61 $\pm$ 0.18 0.69 $\pm$ 0.02 0.69 $\pm$ 0.03 0.69 $\pm$ 0.03
100 -5.79 $\pm$ 7.19 0.66 $\pm$ 0.1 0.64 $\pm$ 0.11 0.68 $\pm$ 0.08 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02
Pendulum 10 2.35 $\pm$ 2.94 -10.76 $\pm$ 7.28 -15.25 $\pm$ 8.86 -9.76 $\pm$ 9.09 4.94 $\pm$ 0.48 5.51 $\pm$ 0.72 5.21 $\pm$ 0.61
25 4.19 $\pm$ 0.89 -0.4 $\pm$ 14.08 0.2 $\pm$ 13.03 4.68 $\pm$ 2.98 5.45 $\pm$ 0.43 6.21 $\pm$ 0.98 6.23 $\pm$ 0.84
50 4.34 $\pm$ 0.75 8.34 $\pm$ 0.83 7.97 $\pm$ 0.74 8.68 $\pm$ 0.72 5.92 $\pm$ 0.32 7.65 $\pm$ 0.5 7.15 $\pm$ 0.76
100 4.69 $\pm$ 0.56 9.53 $\pm$ 0.96 7.07 $\pm$ 5.49 8.8 $\pm$ 1.23 6.09 $\pm$ 0.42 7.66 $\pm$ 0.64 7.77 $\pm$ 0.68
Hopper 10 -31.94 $\pm$ 28.24 -26.65 $\pm$ 15.57 -12.96 $\pm$ 9.55 -10.81 $\pm$ 4.25 9.36 $\pm$ 1.79 8.55 $\pm$ 1.34 8.55 $\pm$ 1.88
25 -5.74 $\pm$ 2.66 12.62 $\pm$ 4.32 13.88 $\pm$ 2.27 17.29 $\pm$ 3.45 13.31 $\pm$ 2.3 13.9 $\pm$ 1.3 13.93 $\pm$ 1.6
50 -2.31 $\pm$ 3.08 27.28 $\pm$ 2.42 23.49 $\pm$ 2.97 28.08 $\pm$ 2.76 16.43 $\pm$ 1.27 17.6 $\pm$ 1.89 16.93 $\pm$ 1.43
100 1.75 $\pm$ 2.83 30.35 $\pm$ 1.44 26.54 $\pm$ 2.19 28.44 $\pm$ 1.52 18.5 $\pm$ 0.65 20.19 $\pm$ 1.37 18.54 $\pm$ 1.23
Ant 10 11.83 $\pm$ 2.29 -4.91 $\pm$ 12.37 3.29 $\pm$ 6.43 10.2 $\pm$ 5.57 13.8 $\pm$ 5.87 22.13 $\pm$ 3.11 23.32 $\pm$ 2.73
25 17.78 $\pm$ 2.28 5.79 $\pm$ 7.68 8.44 $\pm$ 3.68 11.97 $\pm$ 5.41 18.69 $\pm$ 9.06 26.98 $\pm$ 1.64 27.66 $\pm$ 1.78
50 22.21 $\pm$ 2.11 12.24 $\pm$ 3.25 14.94 $\pm$ 3.3 16.16 $\pm$ 3.2 20.54 $\pm$ 8.88 29.84 $\pm$ 1.09 29.62 $\pm$ 0.7
100 23.4 $\pm$ 1.85 15.26 $\pm$ 6.5 16.87 $\pm$ 3.75 19.6 $\pm$ 2.07 20.03 $\pm$ 9.87 31.24 $\pm$ 1.09 31.63 $\pm$ 0.9
Humanoid 10 -230.41 $\pm$ 19.87 -156.59 $\pm$ 82.79 -210.52 $\pm$ 53.23 -242.95 $\pm$ 2.96 -247.83 $\pm$ 1.71 -226.42 $\pm$ 46.74 -194.78 $\pm$ 73.74
25 -245.45 $\pm$ 2.24 -224.95 $\pm$ 37.12 -245.82 $\pm$ 5.03 -247.3 $\pm$ 1.26 -247.83 $\pm$ 1.19 -247.39 $\pm$ 3.18 -232.34 $\pm$ 42.02
50 -246.91 $\pm$ 1.28 -247.38 $\pm$ 3.28 -249.28 $\pm$ 0.52 -249.2 $\pm$ 0.52 -248.06 $\pm$ 1.82 -249.41 $\pm$ 0.73 -223.49 $\pm$ 77.83
100 -247.74 $\pm$ 0.6 -248.97 $\pm$ 1.1 -249.4 $\pm$ 0.6 -249.66 $\pm$ 0.53 -247.25 $\pm$ 3.36 -249.91 $\pm$ 0.12 -223.29 $\pm$ 79.8
@@ Appendix F Introduction to Normalizing Flows
NFs are powerful non-parametric models that have demonstrated the ability to fit flexible multi-modal distributions [54, 53]. These models achieve this by transforming a simple base continuous distribution, such as Gaussian or Beta, into a more complex one using the change of variable formula. By enabling scoring and sampling from the fitted distribution, NFs find application across various problem domains. Let $B$ represent the base distribution, a D-dimensional continuous random vector with $p_{B}(b)$ as its density function, and let $Y=g(B)$, where $g$ is an invertible function with an existing inverse $g^{-1}$, and both $g$ and $g^{-1}$ are differentiable. Leveraging the change of variable formula, we can express the distribution of $Y$ as follows:
@@
where $J(\cdot)$ denotes the Jacobian, and $\det$ signifies the determinant. The first term on the right-hand side of Equation 11 governs the shape of the distribution, while $|\det(J(g^{-1}(y)))|$ normalizes it, ensuring the distribution integrates to one. Complex distributions can be effectively modeled by making $g(b)$ a learnable function with tunable parameters $\theta$, denoted as $g_{\theta}(b)$. However, it is essential to select $g$ carefully to guarantee its invertibility and differentiability. For examples of suitable choices, please refer to [47].
@@ Appendix G Hypothesis Testing Details
We conducted Welch’s t-tests to compare means $(\mu_{i},\mu_{j})$ between different estimators, as this test relaxes the assumption of equal variances compared to other hypothesis tests [13]. The means for both KL and Bhatt were compared to each of the baseline methods: BatchBALD, BADGE, BAIT, MC and random. To control the family-wise error rate (FWER), we performed a Holm-Bonferroni correction across each setting, environment, and acquisition batch. This follows best practices to ensure our results are statistically significant and do not occur just by random chance.